Bukit Barisan (04.01.2010 to 06.30.2010)

You are here

PDF versionPDF version
April - June
Communications activities: 
We have no personnel changes since January. Meyner Nusalawo (Opo) is leading the camera trap and the vegetation protocols.
Protocol Problems: 
Climate protocol: Our climate station was set up and established on 24th February 2010. All systems are running well. We have finally been able to retrieve the data into one file (before the data was separated into daily file system). To retrieve the data into one file, we didn’t use card convert (in PC200W) but use view file option, instead. The 2MB card may not be fit for 1 month data storage. However, we still have problems in retrieving the data every two weeks because of limited staff who is mobile and able to use internet to send the data. Data up to July 13th has been uploaded through the TEAM portal. Camera Trap protocol: Our first and second array camera trap has been deployed April and June 2010. Divided into two teams consisted of four people (2 field staff and 2 porters), four to five days were usually needed to deploy all the camera traps. No particular problems were found during the deployment. Cameras were usually deployed at a range between 0-20m from proposed points depends on potential location for animal paths. First array of camera trap data has been uploaded in deskTEAM 1.0.2 Vegetation protocol: Two vegetation plots have been set up and all trees/lianas have been measured during May 2010. Only one team was used to carry out the vegetation survey. The EcoPDA was tried out but it may need script change.
Schedule Problems: 
Formal presentation to the DirGen of PHKA Dept of Forestry has been delayed for several times since October 2009, and this may affect our schedule since formal approval has not been obtained. However, our climate station has been set up by the end of February. Camera traps from the first and second array have been set up April and June 2010 as well as two vegetations plots that were surveyed on May 2010, a bit late than what was scheduled. This was due to short on staff. Our camera trap leader (Meyner Nusalawo) had to take a leave due to personal matters. In consequence, the rest of the schedule left (4 vegetation plots) need to be adjusted up to early September.
Logistical Problems: 
Climate station has been set up and established, as well as fences around the climate station. However, only one pyranometer was received and this has been set up as well. Several cutting of trees should be carried out around the climate station. WCS Indonesia is in the process of renewing the MOU, and we think that the process of requesting permission to cut trees from the Department of Forestry should be done after the MOU. None other important problems were found during deployment of climate, camera trap, and vegetation protocols. Two camera were not working properly (rapidly flashing red and does not identify the memory card).